OpenStudy (anonymous):

Derivative of 4 sin^2 t?

hartnn (hartnn):

what have u tried?

OpenStudy (anonymous):

I come up with 8 sin^2 t cos^2 t. using the Chain rule

OpenStudy (anonymous):

First things first, with respect to which variable are you derivating?

hartnn (hartnn):

ok, derivative of \(y^2 \) will be 2y dy/dx using chain rule, right ?? so derivative of \(\sin^2 t\) will be \(2\sin t \dfrac{d}{dt} \sin t \) makes sense ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so, whats the derivative of sin t ?

OpenStudy (anonymous):

cos t

hartnn (hartnn):

so whats the final answer :)

OpenStudy (anonymous):

2 sin t cost

OpenStudy (anonymous):

Do i multiply the 4 from the original equation by 2

hartnn (hartnn):

and with the 4 in the beginning, it'll be 8 sin t cost or 3 sin 2t :) oh yes :)

hartnn (hartnn):

i meant 4 sin 2t ****

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

i mean it s 2 sinx cosx

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

How did you arrive at 4 sin 2t?

OpenStudy (anonymous):

@nelsonjedi trigonometric identity \[\sin(2x)=2sinxcosx\]

hartnn (hartnn):

the double angle formula \(\Large 2 \sin x \cos x = \sin 2x \)

OpenStudy (anonymous):

If you have not learned the formula yet, you can leave the answer as \[8\times sint \times cost\]

OpenStudy (anonymous):

However if you are doing differential calculus I assume you to be proficient with your identities....

OpenStudy (anonymous):

Ok i am good with my original solution which I have to square But in checking my answer they show that (8 sin t cost) ^ 2 as being 16 sin t^2 cos t^2...I found it to be 64 sin t^2 cos ^ 2...

OpenStudy (anonymous):

Proficient and remembering are two different things...

hartnn (hartnn):

you have to square your answer? then it'll be \(64 \sin^2 t \cos^2 t\)

OpenStudy (anonymous):

That is what I thought..Thank you again.

hartnn (hartnn):

welcome ^_^